# Practical Global and Local Bounds in Gaussian Process Regression via Chaining

### v1

@@ Results
We illustrate the 99% uncertainty bounds in Figure 3 (Large images are provided in Appendix C.4), which corresponds to the results shown in Table 1. In all figures, our method achieves over 99% coverage (with all black test points within bounds) and consistently produces tighter intervals, indicating its superior performance compared to all baselines. Computational cost results are reported in Appendix C.1. Our method achieves competitive computational efficiency and strong scalability across datasets. Statistical significance results can be found in Appendix C.5. In the vast majority of cases, our method demonstrates statistically significant improvement.
@@ Conclusion
Our work addresses the limitations of existing methods by introducing a novel chaining-based approach that improves error control and robustness. Leveraging Talagrand’s techniques (talagrand2014upper), we derive more flexible and accurate prediction bounds without relying on posterior means or variances. Our method not only yields conventional uncertainty bounds but also estimates the expected value and variance of the supremum using only the training set. Furthermore, it provides tighter bounds for commonly used kernels such as RBF and Matérn. Experiments on synthetic and real-world datasets confirm the performance of our approach.

### v2

@@ Results
We illustrate the 99% uncertainty bounds in Figure 3 (large images are provided in Appendix C.4), which corresponds to the results shown in Table 1. In all figures, our method achieves over 99% coverage (with all black test points within bounds) and consistently produces tighter intervals, indicating its superior performance compared to all baselines. Computational cost results are reported in Appendix C.1. Our method achieves competitive computational efficiency and strong scalability across datasets. Statistical significance results can be found in Appendix C.5. In the vast majority of cases, our method demonstrates statistically significant improvements.
@@ Conclusion
Our work addresses the limitations of existing methods by introducing a novel chaining-based approach that improves error control and robustness. Leveraging Talagrand’s techniques (talagrand2014upper), we derive more flexible and accurate prediction bounds without relying on posterior means or variances. Our method not only yields conventional uncertainty bounds but also estimates the expected value and variance of the supremum using only the training set. Furthermore, it provides tighter bounds for commonly used kernels such as RBF and Matérn. Future work includes case analyses, point-selection strategies, and broader related work.